# Online Conformal Prediction with Efficiency Guarantees

### v1

@@ 4 Lower Bound for Arbitrary Order Sequences
Fix any potentially randomized algorithm $\mathcal{A}$. Consider what $\mathcal{A}$ does on input sequence $S^{(K)}_{K}$. We say that $\mathcal{A}$ “hits” phase $j$ if it captures $\geq\frac{\alpha T}{2}$ points in phase $j$, in expectation. If $\mathcal{A}$ captures $<\frac{\alpha T}{2}$ points in phase $j$ in expectation, we say it “misses.” Either:
@@
    For $S^{(K)}_{j}$, we know that the interval $[0,\varepsilon^{K-j+1}]$ achieves coverage 1 on every phase other than phase $j$. Thus, $\textsc{Opt}_{S^{(K)}_{j}}(\alpha)\leq\varepsilon^{K-j+1}$. This means that on $S^{(K)}_{j}$, $\mathcal{A}$ plays intervals of average volume
@@
$=\Omega\left(\min\left\{\frac{\log(1/\mathsf{minwidth})}{\log(\mu_{\mathrm{avg}})}\alpha^{1+\varepsilon^{\prime}}T,~{}T\right\}\right)$ for any $\text{for any }\varepsilon^{\prime}>0,$
@@
Finally, we note that while this argument considered the performance of $\mathcal{A}$ on intervals with one end at 00 for ease of notation, we could have instead considered a symmetric version of the sequences, where we map each phase that is drawn
@@ 5 Lower Bound Against “Best of Both Worlds” Algorithms
    On every day $t$, $\mathcal{A}$ plays an interval that has expected miscoverage $\geq\frac{1}{2}\cdot\frac{\alpha T}{t}$. Summing over $t$, we see that this leads to $\approx\ln(1/\alpha)\alpha T$ mistakes in expecation overall.
@@
Let $20<\varepsilon\leq 1/$ 2 and $0\leq K\leq\ln(1/\alpha)$ be parameters that we will set later. We define a distribution $\mathcal{D}^{(K)}$ over $[0,1]$, such that
@@
$\operatorname*{\mathbb{P}}_{x\sim\mathcal{D}^{(K)}}\left[x\leq\varepsilon^{1+K-i}\right]$ $=1-\alpha e^{K-i}$ for $\text{for }0\leq i\leq K,$ (15)
$\operatorname*{\mathbb{P}}_{x\sim\mathcal{D}^{(K)}}\left[x\leq\varepsilon+t(1-\varepsilon)\right]$ $=1-\alpha(1-t)$ for $\text{for }0\leq t\leq 1,$
@@
Since the density is non-increasing for $x\geq\varepsilon^{K+1}$, and $\mathcal{D}^{(K)}$ has no mass on $x<\varepsilon^{K+1}$, for any coverage level $0\leq c\leq 1$, a minimum volume set achieving coverage $c$ over $\mathcal{D}^{(K)}$ is of the form $[\varepsilon^{K+1},F_{\mathcal{D}^{(K)}}^{-1}(c)]$, so
@@
Let $1\leq w\leq T$ be a window length that we will set later, and let $W_{t}$ denote the window (sequence of consecutive days) of length $w$ that ends on day $t$. If $\mathcal{A}$ has average expected miscoverage $\leq\frac{1}{e}\frac{\alpha T}{t}$ over $W_{t}$ on input $S^{(K)}$, then we say that $\mathcal{A}$ “hits” $W_{t}$. Otherwise, $\mathcal{A}$ has average expected miscoverage $>\frac{1}{e}\frac{\alpha T}{t}$ over $W_{t}$ on input $S^{(K)}$, and we say that $\mathcal{A}$ “misses” $W_{t}$.
@@ 6 Uniform Convergence for Exchangeable Sequences
First we note from the projection property of exchangeable sequences, the subsequence $\mathbf{S}=(S_{1},\dots S_{t})$ is also exchangeable. The proof of uniform convergence follows the standard symmetrization approach as given in Bousquet, Boucheron, and Lugosi (2004).
@@
$\operatorname*{\mathbb{P}}\Big{[}\ell(h,\mathbf{S})-\ell(h,\mathbf{S}_{t})\geq\lambda\Big{]}\leq\exp\Big{(}-t\lambda^{2}/2\Big{)},\text{ and }\operatorname*{\mathbb{P}}\Big{[}\ell(h,\mathbf{S}^{\prime}_{t})-\ell(h,\mathbf{S})\geq\lambda\Big{]}\leq\exp\Big{(}-t\lambda^{2}/2\Big{)}.$ (20)
@@ Acknowledgements
I thank my advisor Aravindan Vijayaraghavan for helpful discussions, and for the proof that uniform convergence extends to exchangeable sequences. This work was supported by the NSF-funded Institute for Data, Econometrics, Algorithms and Learning (IDEAL) through the grant NSF ECCS-2216970, the NSF via grant CCF-2154100, and the Northwestern Presidential Fellowship.

### v2

@@ 4 Lower Bound for Arbitrary Order Sequences
Fix any potentially randomized algorithm $\mathcal{A}$. Consider what $\mathcal{A}$ does on input sequence $S^{(K)}_{K}$. We say that $\mathcal{A}$ “hits” phase $j$ if it captures $\geq\frac{\alpha T}{2}$ points in phase $j$, in expectation. If $\mathcal{A}$ captures $\frac{\alpha T}{2}$ points in phase $j$ in expectation, we say it “misses.” Either:
@@
    For $S^{(K)}_{j}$, we know that the interval $0,\varepsilon^{K-j+1}]$ achieves coverage 1 on every phase other than phase $j$. Thus, $\textsc{Opt}_{S^{(K)}_{j}}(\alpha)\leq\varepsilon^{K-j+1}$. This means that on $S^{(K)}_{j}$, $\mathcal{A}$ plays intervals of average volume
@@
$=\Omega\left(\min\left\{\frac{\log(1/\mathsf{minwidth})}{\log(\mu_{\mathrm{avg}})}\alpha^{1+\varepsilon^{\prime}}T,~T\right\}\right)$ $\text{for any }\varepsilon^{\prime}>0,$
@@
Finally, we note that while this argument considered the performance of $\mathcal{A}$ on intervals with one end at 0 for ease of notation, we could have instead considered a symmetric version of the sequences, where we map each phase that is drawn
@@ 5 Lower Bound Against “Best of Both Worlds” Algorithms
    On every day $t$, $\mathcal{A}$ plays an interval that has expected miscoverage $\geq\frac{1}{2}\cdot\frac{\alpha T}{t}$. Summing over $t$, we see that this leads to $\approx\ln(1/\alpha)\alpha T$ mistakes in expectation overall.
@@
Let $0<\varepsilon\leq 1/2$ and $0\leq K\leq\ln(1/\alpha)$ be parameters that we will set later. We define a distribution $\mathcal{D}^{(K)}$ over $[0,1]$, such that
@@
$\operatorname*{\mathbb{P}}_{x\sim\mathcal{D}^{(K)}}\left[x\leq\varepsilon^{1+K-i}\right]$ $=1-\alpha e^{K-i}$ $\text{for }0\leq i\leq K,$ (15)
$\operatorname*{\mathbb{P}}_{x\sim\mathcal{D}^{(K)}}\left[x\leq\varepsilon+t(1-\varepsilon)\right]$ $=1-\alpha(1-t)$ $\text{for }0\leq t\leq 1,$
@@
Since the density is non-increasing for $x\geq\varepsilon^{K+1}$, and $\mathcal{D}^{(K)}$ has no mass on $x<\varepsilon^{K+1}$, for any coverage level $0\leq c\leq 1$, a minimum volume set achieving coverage $c$ over $\mathcal{D}^{(K)}$ is of the form $\varepsilon^{K+1},F_{\mathcal{D}^{(K)}}^{-1}(c)]$, so
@@
Let $1\leq w\leq T$ be a window length that we will set later, and let $W_{t}$ denote the window (sequence of consecutive days) of length $w$ that ends on day $t$. If $\mathcal{A}$ has average expected miscoverage $\leq\frac{1}{e}\frac{\alpha T}{t}$ over $W_{t}$ on input $S^{(K)}$, then we say that $\mathcal{A}$ “hits” $W_{t}$. Otherwise, $\mathcal{A}$ has average expected miscoverage $\frac{1}{e}\frac{\alpha T}{t}$ over $W_{t}$ on input $S^{(K)}$, and we say that $\mathcal{A}$ “misses” $W_{t}$.
@@ 6 Uniform Convergence for Exchangeable Sequences
First we note from the projection property of exchangeable sequences, the subsequence $\mathbf{S}=(S_{1},\dots S_{t})$ is also exchangeable. The proof of uniform convergence follows the standard symmetrization approach as given by Bousquet, Boucheron, and Lugosi (2004).
@@
[…]
@@
7 Future Directions
In this work we provide a new framework to measure the efficiency of online conformal prediction algorithms. We apply this to studying the setting where the response variables are in the unit interval $[0,1]$, the notion of efficiency is minimizing the volume (Lebesgue measure) of the prediction set, and the family of sets that we compete against is subintervals $a,b]\subseteq[0,1]$. We study this in the unsupervised (featureless) setting for arbitrary and exchangeable input sequences. We view this as the most basic and fundamental setting in which to explore efficiency optimality. We believe that there are many opportunities for future work to explore efficiency optimality for online conformal prediction in a broad range of other settings that are widely studied in conformal prediction and present new challenges. We provide a few examples of these settings.
  • Other prediction sets. Even in the setting where the target variables live in the unit interval $[0,1]$, there are other families of prediction sets that we can consider competing against. For example, in the standard setting, Gao et al. (2025) provides an algorithm that competes with the volume-optimal union of $k$ intervals. Can we design an algorithm that competes with unions of intervals in the online setting?
  • Other geometries. It would be interesting to extend this result to other spaces of target variables $\mathcal{Y}$. For example, what if instead of being points in the unit interval, our target variables were points in high-dimensional space? Are there natural classes of prediction sets that we can compete against in the online setting?
  • Other notions of efficiency. In this work we consider minimizing the volume (Lebesgue measure) of the prediction sets played. Depending on the task that the conformal predictor is used for, there could be other natural notions of efficiency. For example, we could consider minimizing other measures of the prediction set. Alternatively, the cost of a prediction set could correspond to some downstream decision-making cost. Can we design online conformal prediction algorithms that minimize other natural notions of cost on the sets played?
  • Other forms of distribution shift. In this work, we consider exchangeable sequences, which exhibit no distribution shift, and arbitrary sequences that exhibit arbitrary distribution shifts. One conclusion of this work is that, even in our basic setting, these two types of sequences admit very different guarantees. Are there natural but more benign forms of distribution shift for which we can design algorithms that have stronger guarantees than what is possible for fully arbitrary sequences?
  • Supervised settings and conditional guarantees. In the supervised setting of conformal prediction, each target variable $y_{t}\in\mathcal{Y}$ is associated by a feature variable $x_{t}\in\mathcal{C}$ which may be predictive of it. On each day, the algorithm observes $x_{t}$ and then produces a prediction set $C(x_{t})$ to try to capture $y_{t}$. It would be interesting to explore efficiency optimality in this supervised setting. This could involve interesting new notions of optimality. Now that we have access to the $x_{t}$, it does not necessarily make sense to compete with the best fixed prediction set in hindsight, as that eliminates strategies that vary as a function of $x_{t}$. It would be interesting to use the features to get guarantees against stronger baselines that are allowed to depend on the features.
@@ Acknowledgements
I thank my advisor Aravindan Vijayaraghavan for helpful discussions, and for the proof that uniform convergence extends to exchangeable sequences (Section 6). This work was supported by the NSF-funded Institute for Data, Econometrics, Algorithms and Learning (IDEAL) through the grant NSF ECCS-2216970, the NSF via grant CCF-2154100, and the Northwestern Presidential Fellowship.